-78 is which term of an arithmetic sequence whose common difference is -4 and whose first term is 6?
Answer by stanbon(75887)

-78 is which term of an arithmetic sequence whose common difference is -4 and whose first term is 6?
-----------
a(n) = a(1) + (n-1)d
---
-78 = 6 + (n-1)(-4)
-84 = (n-1)(-4)
n-1 = 21
n = 22
---
The 22nd term is -78
====================
Cheers,
Stan H.